OpenStudy (anonymous):

Evaluate: (17.2)/(25-4^-2)

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{17.2} {25 - 4^{-2}} \)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

keep in mind that \(\bf \Large a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (mathstudent55):

\(\dfrac{17.2} {25 - 4^{-2}}\) \(= \dfrac{17.2} {25 - \frac{1}{4^{2}} } \) \(= \dfrac{17.2} {25 - \frac{1}{16} } \) \(= \dfrac{17.2} {25 - 0.0625 } \)